OpenStudy (superhelp101):

What is the mass, in grams, of a sample of 6.12 × 102^24 atoms of manganese (Mn)?

OpenStudy (superhelp101):

@dhiraj123

OpenStudy (superhelp101):

why did you multiply 54.938

OpenStudy (anonymous):

6.12 × 10^24 atoms(1 mole Mn/6.022 x 10^23 atoms)= 10.16 mole Mn(54.93g/ 1 mole Mn)=558.089g

OpenStudy (anonymous):

6.12 x10^24 atoms (1 mol/6.02 x 10^23)(55 g Mn/ 1 mol Mn) = 559 g

OpenStudy (superhelp101):

oh ok!